# Active operator learning with predictive uncertainty quantification for partial differential equations

### v1

@@ 2.4 Sampling procedure for randomized data generation
To construct the training datasets for the DeepONet models, we first employ Gaussian process sampling [44] to generate random realizations of the solution functions. In particular, we sample
@@
from mean-zero Gaussian random fields with squared exponential covariance kernels $k_{l}(x,y)\,=\,\exp(-\|x-y\|^{2}/2l^{2})$. Training data was generated using the length-scales $l\in\{0.2000,0.2333,0.2667,$ $0$, and additional test data was generated for length-scales $l\in\{0.1333,0.1667,0.3333,0.3667\}$. The test length-scales were intentionally selected outside of the range used to generate training data so that we can assess how well the proposed models generalize to out-of-distribution data.
@@ 3 Validation of the proposed framework
The first case corresponds to the network generalization on in-distribution data, while the latter corresponds to the generalization for out-of-distribution (OoD) data [45]. We first demonstrate that the network predictions are unbiased, and produce accurate predictions, on a set of linear problems. This is followed by a detailed analysis of the predictive uncertainty estimates. We then assess the network architecture’s generalization capabilities on two nonlinear problem setups.
@@ 3.1 Poisson Equations with Homogeneous and Inhomogeneous BCs
The first problem we consider is the Poisson equation on a square domain with homogeneous boundary conditions, as described by Eq. (2a). This is a simple problem, and if the proposed model performs poorly in this setting, we can conclude that the proposed addition of UQ is negatively impacting accuracy. Overall, the network predictions for this problem setup are accurate, as shown in Table 1, with relative errors below $1\%$. This allows us to move on to more complex problems with confidence that the UQ framework is not undermining the accuracy of network predictions.
@@ 3.2 Assessment of predictive uncertainty framework
The biases of the error distributions at the tail-end of training are consistently small, with values ranging from $-0.0018$ to $0.0019$. The histogram of biases is also symmetric and roughly centered at zero, indicating that the network does not have any tendency to over-estimate or under-estimate function values consistently across the dataset. The skew values for the observed error distributions are also approximately symmetric and centered at zero, with a minimum skew of $-0.4739$ and maximum skew of $0.3715$. The majority of error distributions have skews with magnitudes below $0.2$, indicating the distributions are roughly symmetric. Since we elected to use normal distributions to model uncertainty, it is important to verify the observed error distributions are sufficiently symmetric. If the observed errors have skewed distributions, the loss function in Eq. (1) can be replaced with the negative log-likelihood for a different class of distributions (see Appendix B of [36]).
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 6: Biases (left), skews (center-left), and kurtosis (center-right) of the observed error distributions at the tail-end of training for $500$ validation examples from the Inhomogeneous Poisson equation over a circular domain. To the right, we show the observed standard deviations in errors at the end of training (blue) and the network’s predicted uncertainties (pink) for 500 examples from the Inhomogeneous Poisson; the significant overlap (maroon) provides evidence that the network uncertainty is accurately capturing the model errors.
A similar analysis of the kurtosis of the observed errors shows a generally positive and skewed right distribution. Kurtosis measures the “tailedness” of a distribution, with a higher value indicating more outliers [46]. Most of the kurtosis values are close to zero with a slight bias in the positive direction. While this means they will generally have slightly longer tails and sharper peaks, all kurtosis values are well within the standard reference interval, $[-2,2]$, for a normal distribution.
@@ 3.4 Generalization to Out-of-Distribution Data
To assess the proposed uncertainty framework’s generalization capabilities, we evaluate the model’s performance on unseen in-distribution and out-of-distribution data. In the presence of sufficient training data, the prediction accuracy of the DeepONet models on in-distribution validation data matches the accuracy on training examples across all length-scales. The associated uncertainty estimates are also well-calibrated to the validation errors when sufficient training data is available, as shown in Fig. 8.
Figure 8: Comparison of DeepONet performance on training and validation examples for different training dataset sizes for the Nonlinear Diffusion-Reaction problem setup (left). The network converges well even with limited training data, but the performance is also observed to level off quickly as the number of training examples increased indicating data saturation and a possible limit to the network’s accuracy. To the right, we provide an analysis of predictive uncertainty produced by a DeepONet model on the Nonlinear Diffusion-Reaction problem setup. When evaluated on in-distribution validation data (blue) as well as out-of-distribution test data (orange) the network’s uncertainty is seen to accurately capture the observed model error.
@@
This shows that the network can, to some extent, automatically determine the difficulty associated with a given problem. More precisely, the accuracy of the network’s predictive uncertainties extends beyond the scope of the function classes observed during training. The network is never provided any labeling information regarding the length-scale values, so the network’s predictive uncertainty is based solely on features extracted from the raw input function values.
4 Accelerating active learning with predictive uncertainties
In this section, we demonstrate how the proposed uncertainty framework can be used to accelerate outer-loop active learning procedures. We consider a system of time-dependent functions parameterized by a vector $\theta=(\alpha,\beta)$ with $\alpha\in[-7,4]$ and $\beta\in[0,6]$. The system consists of an acceleration $f_{\theta}(t)$, velocity $u_{\theta}(t)$, and final position $q_{\theta}$ defined by
@@ 4 Accelerating active learning with predictive uncertainties
[…]
@@
[…]
@@
We will identify the optimal velocity function using as few system queries as possible; this is a common requirement for practical applications where data collection is expensive and/or time-consuming and QoI estimates must be made using a limited budget of system interrogations. For improved data efficiency, we employ active learning to iteratively refine the DeepONet model and direct our search for the optimal system configuration. The governing equations for this problem were specifically designed to include two peaks inside the constrained parameter space: a local maximum at $(-4.7680,\pi)$ and a global maximum at $(1.2150,\pi)$, as shown in Fig. 9. We start with a random selection of points with $\alpha<-2.0$, so that the initial observations are confined to the left half of the domain with a bias toward the local maximum.
@@
The proposed active learning procedure leverages the uncertainty estimates of the DeepONet models to explore the parameter space in regions of high uncertainty while also exploiting the mean predictions to search regions with high predicted QoI values. To assess the performance of the active learning procedure, we compare its results with greedy (exploitation-driven) and random (exploration-driven) alternatives. Details for the three strategies are summarized below:
  • Active Learning with UQ: New query locations are proposed based on the upper confidence bound induced by the network’s predictive uncertainty. We evaluate the network on a grid of parameter values, $\Lambda\subset[-7,4]\times[0,6]$, and integrate the network’s velocity predictions to obtain mean and standard deviation estimates for the QoI. We use the acquisition values $\widehat{q}_{\theta}+\varepsilon\cdot\sigma(\widehat{q}_{\theta})$ to rank the query locations $\theta\in\Lambda$, where $\varepsilon>0$ isused to weight the impact of uncertainty. The system is then evaluated at the locations with the highest acquisition values, and the network is re-trained with the augmented dataset before the next active learning iteration.
@@
For both the active learning and greedy strategies, a degree of random exploration is introduced to balance exploitation and exploration. During the initial iterations, we supplement the model’s proposed query locations with additional points sampled uniformly from the parameter grid $\Lambda$. We found this significantly improved the consistency of both approaches and helped prevent instances of local trapping [^3] Both the active and greedy strategies occasionally converged to the local maximum during our initial experiments. This problem was resolved by incorporating random exploration to supplement both search strategies. .
@@ 4.1 Parameter-to-Function Learning
In the first formulation, we feed the parameters $(\alpha,\beta)$ directly into the branch network of the DeepONet and train the trunk network using randomly sampled time locations on the interval $[0,1]$. An initial training dataset is created by randomly sampling $50$ parameter locations in $[-7,-2]\times[0,6]$ and computing the associated velocity functions. The DeepONet model is tasked with approximating the mapping $\theta\mapsto u_{\theta}(t)$, and we use the predictive uncertainty estimates to compute the value of the upper confidence bound acquisition function [47]. This acquisition function is used to select the next batch of queries for each active learning iteration. To obtain batches of distinct query locations, we define a selection distance, $\delta_{\tau}>0$, that specifies the minimum allowable distance between queries proposed at the current iteration, $\tau$. We achieve this by selecting query locations sequentially according to the following rule:
@@
At each learning iteration, we evaluate the system at $10$ query locations selected using Eq. (4) for the active setup, the highest predicted QoI for the greedy strategy, and uniformly sampled locations for the random approach. To assist with exploration, we also include $5$ randomly sampled query locations during the initial iteration. We reduce the number of random points added by a factor of $0.9$ for subsequent iterations, rounding decimal values to the nearest integer. We perform $20$ learning iterations for each strategy, resulting in a total of $294$ system queries. We conducted $50$ independent trials for each strategy to assess performance across different initial sample locations. The initial data locations were sampled randomly for each of the $50$ trials, and each learning strategy used the same initial samples to provide a fair comparison of the results.
@@
Figure 10: Results for parameter-to-function active learning problem. Plots of the final predicted position field at time $t=1$ (top left), the absolute error relative to the ground truth (top center), and the model’s predicted uncertainty in position (top right). By construction of the active learning setup, the model is most accurate near the peak location, $(1.2150,\pi)$. The uncertainty in position is derived from the model’s UQ predictions for velocity and give a reasonable approximation of the observed errors. We conducted $50$ trials for each learning strategy, with each trial consisting of $20$ active learning iterations. The median and spread between the $0.2$-quantile and $0.8$-quantile are plotted (bottom) for each iteration to summarize the results across trials. The errors in the predicted maximum function value are shown on the left, and errors in the predicted maximal location are shown to the right.
@@
Examples of ground-truth and predicted velocity functions are shown in Fig. 9, along with the network’s predicted uncertainty bounds. These mean and uncertainty functions are integrated over time to obtain estimates for the final position (and uncertainty in position). A qualitative summary of the position predictions produced by the active learning procedure are shown in Fig. 10. We note that the error is lowest in a neighborhood around the global peak, indicating the active learning procedure correctly targeted this region for refinement. The uncertainty estimates also align relatively well with the observed errors; however, the network uncertainty underestimates errors in the region $\alpha<0$, possibly due to the limited amount of training data available.
@@
We assess the performance of each learning strategy using two metrics: the error in the predicted maximum QoI value and the error in the maximizing parameters for the QoI. The ground truth solution for the maximum value is $q_{\theta^{*}}\approx 1.3410$, which occurs at $\theta^{*}=(1.2150,\pi)$. A quantitative comparison between the active, greedy, and random strategies is provided in Fig. 10, which summarizes the results from $50$ trials of each learning strategy. The uncertainty-equipped DeepONet model achieves the lowest errors in both the maximum value and maximizing parameter estimates. For the maximum value metric, the active setup achieves a median final error of $48.96\mathrm{e-}$ 4 over $50$ trials, while the greedy and random strategies had median errors of $36.90\mathrm{e-}$ 3 and $25.59\mathrm{e-}$ 2, respectively. The performance difference for parameter estimation was less pronounced, but the active setup still achieved the lowest median error, $0.067$, compared with median errors of $0.133$ and $0.219$ for the greedy and random approaches.
@@ 4.2 Function-to-Function Learning
For the second experiment, we modify the branch input of the DeepONet to use time-series observations of acceleration instead of the parameters $\alpha$ and $\beta$. We create an initial training dataset by randomly sampling $10$ parameter locations in the interval $[-7,-2]\times[0,6]$ and computing the associated acceleration and velocity functions. The DeepONet model is then tasked with predicting velocity functions based on time-series information regarding acceleration.
Initial experiments showed the direct function-to-function mapping was too simple to benefit from active learning. The mapping from acceleration to velocity amounts to a straight-forward integral estimate, and the models provided accurate approximations with minimal system interrogations. To increase the problem complexity, we modify the acceleration information passed to the network using the following transformation:
[…]

### v3

@@ 2.4 Sampling procedure for randomized data generation
To construct the training datasets for the DeepONet models, we first employ Gaussian process sampling Rasmussen [2003] to generate random realizations of the solution functions. In particular, we sample
@@
from mean-zero Gaussian random fields with squared exponential covariance kernels $k_{l}(x,y)\,=\,\exp(-\|x-y\|^{2}/2l^{2})$. Training data was generated using the length-scales $l\in\{0.2000,0.2333,0.2667,$ 0.3000}0.3000\}, and additional test data was generated for length-scales $l\in\{0.1333,0.1667,0.3333,0.3667\}$. The test length-scales were intentionally selected outside of the range used to generate training data so that we can assess how well the proposed models generalize to out-of-distribution data.
@@ 3 Validation of the proposed framework
The first case corresponds to the network generalization on in-distribution data, while the latter corresponds to the generalization for out-of-distribution (OoD) data Hendrycks and Gimpel [2016]. We first demonstrate that the network predictions are unbiased, and produce accurate predictions, on a set of linear problems. This is followed by a detailed analysis of the predictive uncertainty estimates. We then assess the network architecture’s generalization capabilities on two nonlinear problem setups.
@@ 3.1 Poisson Equations with Homogeneous and Inhomogeneous BCs
The first problem we consider is the Poisson equation on a square domain with homogeneous boundary conditions, as described by Eq. (2a). This is a simple problem, and if the proposed model performs poorly in this setting, we can conclude that the proposed addition of UQ is negatively impacting accuracy. The models were trained using the Adam optimization algorithm Kingma and Ba [2014], with a learning rate of $7.5e\mathrm{-}4$, a batch size of $512$, and exponential learning rate decay by a factor of $0.95$ every $25$ thousand learning steps. Training was performed for $2$ million steps, and this setup was used to train all models evaluated in this section. Overall, the network predictions for this problem setup are accurate, as shown in Table 1, with relative errors below $1\%$. This allows us to move on to more complex problems with confidence that the UQ framework is not undermining the accuracy of network predictions.
@@ 3.2 Assessment of predictive uncertainty framework
The biases of the error distributions at the tail-end of training are consistently small, with values ranging from 0.0018-0.0018 to $0.0019$. The histogram of biases is also symmetric and roughly centered at zero, indicating that the network does not have any tendency to over-estimate or under-estimate function values consistently across the dataset. The skew values for the observed error distributions are also approximately symmetric and centered at zero, with a minimum skew of 0.4739-0.4739 and maximum skew of $0.3715$. The majority of error distributions have skews with magnitudes below $0.2$, indicating the distributions are roughly symmetric. Since we elected to use normal distributions to model uncertainty, it is important to verify the observed error distributions are sufficiently symmetric. If the observed errors have skewed distributions, the loss function in Eq. (1) can be replaced with the negative log-likelihood for a different class of distributions (see Appendix B of Winovich et al. [2019]).
Figure 6: Biases (top-left), skews (top-center), and kurtosis (top-right) of the observed error distributions at the tail-end of training for $500$ validation examples from the Inhomogeneous Poisson equation over a circular domain. We show the observed standard deviations (bottom-left) in errors at the end of training (blue) and the network’s predicted uncertainties (pink) for 500 examples from the Inhomogeneous Poisson; the significant overlap (maroon) provides evidence that the network uncertainty is accurately capturing the model errors. There is strong agreement between the true and predicted standard deviations with an $R^{2}$ of 0.977 (bottom-right).
A similar analysis of the kurtosis of the observed errors shows a generally positive and skewed right distribution. Kurtosis measures the “tailedness” of a distribution, with a higher value indicating more outliers Kline [2023]. Most of the kurtosis values are close to zero with a slight bias in the positive direction. While this means they will generally have slightly longer tails and sharper peaks, all kurtosis values are well within the standard reference interval, [2,2][-2,2], for a normal distribution.
@@ 3.4 Generalization to Out-of-Distribution Data
To assess the proposed uncertainty framework’s generalization capabilities, we evaluate the model’s performance on unseen in-distribution and OoD data. In the presence of sufficient training data, the prediction accuracy of the DeepONet models on in-distribution validation data matches the accuracy on training examples across all length-scales. The associated uncertainty estimates are also well-calibrated to the validation errors when sufficient training data is available, as shown in Fig. 8.
Figure 8: Comparison of DeepONet performance on training and validation examples for different training dataset sizes for the Nonlinear Diffusion-Reaction problem setup (left). The network converges well even with limited training data, but the performance is also observed to level off quickly as the number of training examples increased indicating data saturation and a possible limit to the network’s accuracy. To the right, we provide an analysis of predictive uncertainty produced by a DeepONet model on the Nonlinear Diffusion-Reaction problem setup. When evaluated on in-distribution validation data (blue) as well as OoD test data (orange) the network’s uncertainty is seen to accurately capture the observed model error.
@@
This shows that the network can, to some extent, automatically determine the difficulty associated with a given problem. More precisely, the accuracy of the network’s predictive uncertainties extends beyond the scope of the function classes observed during training. The network is never provided any labeling information regarding the length-scale values, so the network’s predictive uncertainty is based solely on features extracted from the raw input function values. This indicates that, although the network is trained under a deterministic regime without explicit epistemic modeling, its predictive uncertainties exhibit epistemic-like behavior by responding to out-of-distribution inputs beyond the training function classes.
4 Accelerating Bayesian optimization with predictive uncertainties
In this section, we demonstrate how the proposed uncertainty framework can be used to accelerate outer-loop BO procedures. As a proof-of-concept problem, we consider a visually intuitive nonlinear system of time-dependent functions parameterized by a vector $\theta=(\alpha,\beta)$ with $\alpha\in[-7,4]$ and $\beta\in[0,6]$. The system consists of an acceleration $f_{\theta}(t)$, velocity $u_{\theta}(t)$, and final position $q_{\theta}$ defined by
@@ 4 Accelerating Bayesian optimization with predictive uncertainties
$q_{\theta}\,$ $=$ $\int_{0}^{1}u_{\theta}(t)\,dt\,\,\,=\,\,\frac{1+\alpha}{\alpha}\,\sin(\alpha)\,-\,e^{\cos(\beta)}.$
@@
$\widehat{q}_{\theta}\,\,=\,\,\int_{0}^{1}\widehat{u}(t)\,dt\,\,\,\approx\,\,\,\frac{1}{2n}\left(\widehat{u}(t_{0})\,+\,\widehat{u}(t_{n})\,+\,2\,\sum\nolimits_{i=1}^{n-1}\,\widehat{u}(t_{i})\,\right)$
@@
We will identify the optimal velocity function using as few system queries as possible; this is a common requirement for practical applications where data collection is expensive and/or time-consuming and QoI estimates must be made using a limited budget of system interrogations. For improved data efficiency, we employ BO to iteratively refine the DeepONet model and direct our search for the optimal system configuration. The governing equations for this problem were specifically designed to include two peaks inside the constrained parameter space: a local maximum at $(-4.7680,\pi)$ and a global maximum at $(1.2150,\pi)$, as shown in Fig. 9. We start with a random selection of points with $\alpha<-2.0$, so that the initial observations are confined to the left half of the domain with a bias toward the local maximum.
@@
The proposed BO procedure leverages the uncertainty estimates of the DeepONet models to explore the parameter space in regions of high uncertainty while also exploiting the mean predictions to search regions with high predicted QoI values. To assess the performance of the BO procedure, we compare its results with greedy (exploitation-driven) and random (exploration-driven) alternatives. Details for the three strategies are summarized below:
  • Bayesian optimization with UQ: New query locations are proposed based on the upper confidence bound induced by the network’s predictive uncertainty. We evaluate the network on a grid of parameter values, $\Lambda\subset[-7,4]\times[0,6]$, and integrate the network’s velocity predictions to obtain mean and standard deviation estimates for the QoI. We use the acquisition values $\widehat{q}_{\theta}+\varepsilon\cdot\sigma(\widehat{q}_{\theta})$ to rank the query locations $\theta\in\Lambda$, where $\varepsilon>0$ isused to weight the impact of uncertainty. The system is then evaluated at the locations with the highest acquisition values, and the network is re-trained with the augmented dataset before the next iteration.
@@
For both the BO and greedy strategies, a degree of random exploration is introduced to balance exploitation and exploration. During the initial iterations, we supplement the model’s proposed query locations with additional points sampled uniformly from the parameter grid $\Lambda$. We found this significantly improved the consistency of both approaches and helped prevent instances of local trapping [^3] Both the BO and greedy strategies occasionally converged to the local maximum during our initial experiments. This problem was resolved by incorporating random exploration to supplement both search strategies. . We provide detailed summaries of the network architectures and training procedures in B.
@@ 4.1 Parameter-to-Function Learning
In the first formulation, we feed the parameters $(\alpha,\beta)$ directly into the branch network of the DeepONet and train the trunk network using randomly sampled time locations on the interval $[0,1]$. An initial training dataset is created by randomly sampling $50$ parameter locations in $7,-2]\times[0,6]$ and computing the associated velocity functions. The DeepONet model is tasked with approximating the mapping $\theta\mapsto u_{\theta}(t)$, and we use the predictive uncertainty estimates to compute the value of the upper confidence bound acquisition function Cox and John [1992]. This acquisition function is used to select the next batch of queries for each learning iteration. To obtain batches of distinct query locations, we define a selection distance, $\delta_{\tau}>0$, that specifies the minimum allowable distance between queries proposed at the current iteration, $\tau$. We achieve this by selecting query locations sequentially according to the following rule:
@@
At each learning iteration, we evaluate the system at $10$ query locations selected using Eq. (4) for the BO setup, the highest predicted QoI for the greedy strategy, and uniformly sampled locations for the random approach. To assist with exploration, we also include $5$ randomly sampled query locations during the initial iteration. We reduce the number of random points added by a factor of $0.9$ for subsequent iterations, rounding decimal values to the nearest integer. We perform $20$ learning iterations for each strategy, resulting in a total of $294$ system queries. We conducted $50$ independent trials for each strategy to assess performance across different initial sample locations. The initial data locations were sampled randomly for each of the $50$ trials, and each learning strategy used the same initial samples to provide a fair comparison of the results.
@@
Figure 10: Results for parameter-to-function BO problem. Plots of the final predicted position field at time $t=1$ (top left), the absolute error relative to the ground truth (top center), and the model’s predicted uncertainty in position (top right). By construction of the BO setup, the model is most accurate near the peak location, $(1.2150,\pi)$. The uncertainty in position is derived from the model’s UQ predictions for velocity and give a reasonable approximation of the observed errors. We conducted $50$ trials for each learning strategy, with each trial consisting of 2020 learning iterations. The median and spread between the $0.2$-quantile and $0.8$-quantile are plotted (bottom) for each iteration to summarize the results across trials. The errors in the predicted maximum function value are shown on the left, and errors in the predicted maximal location are shown to the right.
@@
Examples of ground-truth and predicted velocity functions are shown in Fig. 9, along with the network’s predicted uncertainty bounds. These mean and uncertainty functions are integrated over time to obtain estimates for the final position (and uncertainty in position). A qualitative summary of the position predictions produced by the BO procedure are shown in Fig. 10. We note that the error is lowest in a neighborhood around the global peak, indicating the learning procedure correctly targeted this region for refinement. The uncertainty estimates also align relatively well with the observed errors; however, the network uncertainty underestimates errors in the region $\alpha<0$, possibly due to the limited amount of training data available.
@@
We assess the performance of each learning strategy using two metrics: the error in the predicted maximum QoI value and the error in the maximizing parameters for the QoI. The ground truth solution for the maximum value is $q_{\theta^{*}}\approx 1.3410$, which occurs at $\theta^{*}=(1.2150,\pi)$. A quantitative comparison between the BO, greedy, and random strategies is provided in Fig. 10, which summarizes the results from $50$ trials of each learning strategy. The uncertainty-equipped DeepONet model achieves the lowest errors in both the maximum value and maximizing parameter estimates. For the maximum value metric, the BO setup achieves a median final error of $8.96\mathrm{e-}4$ over $50$ trials, while the greedy and random strategies had median errors of $6.90\mathrm{e-}3$ and $5.59\mathrm{e-}2$, respectively. The performance difference for parameter estimation was less pronounced, but the BO setup still achieved the lowest median error, $0.067$, compared with median errors of $0.133$ and $0.219$ for the greedy and random approaches.
@@ 4.2 Function-to-Function Learning
For the second experiment, we modify the branch input of the DeepONet to use time-series observations of acceleration instead of the parameters $\alpha$ and $\beta$. We create an initial training dataset by randomly sampling $10$ parameter locations in the interval $7,-2]\times[0,6]$ and computing the associated acceleration and velocity functions. The DeepONet model is then tasked with predicting velocity functions based on time-series information regarding acceleration.
Initial experiments showed the direct function-to-function mapping was too simple to benefit from BO. The mapping from acceleration to velocity amounts to a straight-forward integral estimate, and the models provided accurate approximations with minimal system interrogations. To increase the problem complexity, we modify the acceleration information passed to the network using the following transformation:
$f_{\theta}^{input}(t)\,\,=\,\,\left(\,1\,+\,\exp(-\tfrac{1}{2}f_{\theta}(t))\,\right)^{-1}\,-\,\tfrac{1}{2}.$
@@
Figure 11: Results for function-to-function BO problem. The final predicted position field at time $t=1$ for the function-to-function setup (top left). The predicted positions accurately reflect the ground truth values everywhere except the far-right side of the domain, as seen in the plot of observed errors (top center), which is consistent with the predicted uncertainty (top right). We conducted $50$ trials for each learning strategy, with each trial consisting of 2020 learning iterations. The median and spread between the $0.2$-quantile and $0.8$-quantile are plotted for each iteration to summarize the results across trials. The error in the predicted optimal value (bottom left) and predicted optimal location (bottom right) are plotted for each learning strategy. The BO strategy enabled by predictive uncertainty estimates yields an order of magnitude improvement in optimal value prediction.